OpenStudy (anonymous):

How do I solve 72+4-14d=36

OpenStudy (anonymous):

72+4-14d=36 76-14d=36 -14d= -40 d= -40/-14 = 20/7

OpenStudy (anonymous):

Use the order of operations!!! They are 1. parenthesis 2.exponents 3.mult. and div. 4.add and subtract So...... what you should do is write it all out in sections by the equations that have answers and work out the problem bit by bit. It will take awhile but really if you think about it, its easy once you break it down in pieces. REMEMBER!!! use the order of operations with the chunks. Go slowly and check your work.

OpenStudy (anonymous):

Hope this helps!!!!